The rate constant, k, of a certain reaction can be determined using the following formula, k=Cx^2Cy^3/Cz^2Cw^5, where w, x, y and z are the concentrations of four different chemicals.
Find Cy.
Cy =

Answer by Boreal(15235)

I'm assuming Cy=[y], and C is not a variable but the symbol for concentration.
k*z^2*w^5/x^2=y^3
y=k*z^2*w^5/x^2
y= k*w*(z*w/x)^(2/3)=k*Cw(Cz*Cw/Cx)^(2/3)